the army gives a battery of exams to all new recruits. one exam measures a person's ability to work with technical machinery. this exam was given to a random sample of 360 new recruits. in the table below x-scores on the test and f is the frequency of new recruits with score x.

x 1 2 3 4    5 6 7    8    9    10

f    28    42 79 83    51 36 18    12    7    4

a) find the probability distribution for these scores

b) draw a histogram of the probability distribution

c). the ground to air missile battalion needs people with a score of 7 or higher on the exam. What is the probability that a new recruit will meet this criterion?

d). the kitchen battalion can use people with scores of three or less. What is the probability that a new recruit is not overqualified for this work?

E). compute the expected value of these scores

f). compute the standard deviation of these scores
Answer #1

a) Let X denotes the scores on the test.

The probability distribution of X is

X Frequency P(X=x) = ( Frequency / Total frequency) x* P(X=x) x2* P(X=x)
1 28 0.0778 0.0778 0.0778
2 42 0.1167 0.2333 0.4667
3 79 0.2194 0.6583 1.9750
4 83 0.2306 0.9222 3.6889
5 51 0.1417 0.7083 3.5417
6 36 0.1000 0.6000 3.6000
7 18 0.0500 0.3500 2.4500
8 12 0.0333 0.2667 2.1333
9 7 0.0194 0.1750 1.5750
10 4 0.0111 0.1111 1.1111
Total 360 1.0000 4.1028 20.6194

b) Histogram

Using R

> x=1:10
> f=c(28,42,79,83,51,36,18,12,7,4)
> y=rep(x,f)
> b=seq(0.5,10.5,1)
> hist(y,breaks=b,xlab="X", ylab="frequency",main ="Histogram")
frequency 0 20 40 60 80 Histogram

c) Required probability = P ( X> =7)

= P ( X=7) + P( X=8) + P(X=9) + P(X=10)

= 0.0500 + 0.0333 + 0.0194 + 0.0111

= 0.1139

d) Required probability = P ( X < = 3)

= P ( X=1) + P(X=2) + P(X=3)

= 0.0778 + 0.1167 + 0.2194

= 0.4139

e) Expected value of scores

E(X) =sum ( x * P(X=x)) = 4.1028

f) Var(X) = E(X2) - (E(X)) 2

E(X2) = sum ( x2 * P(X=x)) =20.6194

Var(X) = 20.6194 - 4.10282 = 3.7867

Standard deviation of the scores is

SD(X) = sqrt(Var(X)) = 1.9459.